Y=2x+1. Show how to solve for x and for y. Also how to find x and y intercepts with domains of 0,1,2,3.
Answer by rfer(16322)

domain=x
range=y
y=2(0)+1
y=1
(0,1)y int is when x=0
-----------
y=2*1+1
y=3
(1,3)
-------------
y=2*2+1
y=5
(2,5)
-----------------
y=2x+1
0=2x+1
2x=-1
x=-1/2
(-1/2,0) x int is when y = 0